Consider the insoluble compound nickel(II) carbonate, NiCO3. The nickel ion also forms a complex with cyanide ions. Write a nA solution contains 1.80x10-2M lead nitrate and 1.80x10-2M magnesium acetate. Solid sodium fluoride is added slowly to this mA solution contains 0.14 M potassium sulfide and 0.14 M sodium sulfite. Solid silver acetate is added slowly to this mixture.

please help i cant figure them out
Answer #1

Question 1.

NiCO3(s) + 4CN-(aq) <----> [Ni(CN)4]2-(aq) + CO32-(aq)

K = Ksp * Kf = 1.42*10-7 * 2*1031 = 2.84*1024

Explanation: The above equation can be obtained by the adding the two equilibria mentioned below.

NiCO3(s) <----> Ni2+(aq) + CO32-(aq); Ksp = 1.42*10-7 ---- Equilibrium 1

Ni2+(aq) + 4CN-(aq) <----> [Ni(CN)4]2-(aq); Kf = 2*1031---- Equilibrium 2